The length of a rectangle is 2.1cm more than 4 times the width. If the perimeter of the rectangle is 70.2cm, what are its dimensions?

Answer by addingup(3677)

L= 4W+2.1
2L+2W= 70.2 Substitute for L
2(4W+2.1)+2W= 70.2
8W+4.2+2W= 70.2
10W= 66
W= 6.6 And L:
4W+2.1= 4(6.6)+2.1= 28.5
----------------------------------
Check:
2(28.5)+2(6.6)= 70.2
57+13.2= 70.2
70.2= 70.2 We have the correct answer.